How does the following affect the measures of Y, C, and NX (#’s please)? : You purchased a good (Virginia) for $20 that the store had acquired a week earlier (i.e., in that same period) from a producer in Mexico for $12 (worth of pesos).
Answer #1

Answer:

As we know

Y=C+I+G+Nx

Since we have purchased the good at $20 so C will increase by $20

and since the store acquired this good from maxico at $2 so Nx will decrease by $12

Since all investment and government spending is constant so Y will increase by Y=20-12=$8